In the interval 0°C to 660°C plantinum resistance thermometer is used for interpolating temperature on internation scale. If the variation of resistance with temperature is given by R1=R°[1+at+bt2],find the value of a,b given R°=10π,R100=13.948π and R444.6=24.817π

1
Expert's answer
2022-03-04T08:31:03-0500

"1.3948=1+100a+100^2b),"

"2.4817=(1+444.6a+444.6^2b),"

"\implies"

"a=4.13\cdot 10^{-3},"

"b=-1.79\cdot 10^{-6}."